### v1

@@ Methods
No Poverty (SDG-1): SustainFM includes an asset wealth index regression task to assess localised poverty using Demographic and Health Surveys (DHS) data from SustainBench [36]. Wealth is estimated at the cluster level using a principal component analysis (PCA) of household assets, yielding a scalar index averaged across surveyed households in each cluster. The dataset comprises 2,079,036 households across 86,936 clusters in 48 countries, based on DHS surveys conducted between 1996 and 2019.
@@
Zero Hunger (SDG-2): SustainFM includes the Cropland Change Detection (CLCD) data set to support food security and sustainable agricultural land management [37]. CLCD comprises 600 paired satellite image samples (512×512 pixels) from Gaofen-2, captured in Guangdong Province, China, between 2017 and 2019. Each pair is annotated with binary labels that indicate cropland conversion, including changes to buildings, roads, lakes, and bare soil, key threats to agricultural sustainability.
Good Health and Wellbeing (SDG-3): SustainFM incorporates a children’s health regression task using DHS survey data from SustainBench [36]. Children’s health is measured via cluster-level child mortality rates, defined as the number of deaths per 1,000 children under age five within a defined spatial unit. The dataset comprises mortality rates for 105,582 clusters, derived from 1,936,904 children surveyed between 1996 and 2019.
@@
Quality Education (SDG-4): SustainFM includes an analysis of women’s educational attainment, based on the DHS survey dataset from SustainBench [36]. The educational attainment metric is derived by calculating the cluster-level mean of education years for women aged 15-49. Education years are capped at 18 to mitigate outliers, a standard survey threshold. The dataset includes 122,435 cluster-level labels, computed from data on 3,013,286 women across 56 countries.
Gender Equality (SDG-5): SustainFM includes a women’s body mass index (BMI) regression task, using satellite imagery and survey data from SustainBench [36]. This dataset includes 94,866 cluster-level BMI labels, derived from 1,781,403 women of reproductive age (15–49), excluding pregnant individuals.
@@
Clean Water (SDG-6): SustainFM includes an urban black and odorous water (BOW) detection task using the RSBD dataset [38], which comprises 329 multispectral images from the Gaofen-2 satellite in Yantai, China. This region has faced escalating pollution and freshwater shortages. Pixel-level annotations of BOW facilitate the training of segmentation models to detect contaminated water and support water quality management in urban settings.
@@
Affordable and Clean Energy (SDG-7): SustainFM includes a solar farm identification task based on the GloSoFarID dataset [39], which provides 13,703 multispectral Sentinel-2 image samples with pixel-level annotations. The dataset includes global solar farm samples from 2017–2018 and U.S. samples from 2022–2023, covering installations at 10m spatial resolution. These EO-derived maps support the analysis of solar infrastructure expansion and guide decision-making in clean energy planning.
Decent Work and Economic Growth (SDG-8): SustainFM includes a task for detecting settlements without electricity using the DSE dataset [40], developed for the 2021 IEEE GRSS Data Fusion Contest. The dataset spans Malawi’s Dezda and Salima districts, covering 98 EO tiles of 800×800 pixels each ( 64 km²). It combines Sentinel-1, Sentinel-2, Landsat 8, and VIIRS nighttime imagery, labeling areas based on both residential presence and electricity access across four categorical classes. This task supports monitoring of energy poverty and associated economic disparities.
Industry, Innovation, and Infrastructure (SDG-9): SustainFM includes a smoke plume detection task based on the SmokePlume dataset [41], which supports model development for industrial impact assessment and environmental monitoring. The dataset builds on 624 industrial sites listed in the European Pollutant Release and Transfer Register, with Sentinel-2 Level-2A imagery from 2019. It includes 21,350 multispectral image patches (120 × 120 pixels; 1.2 × 1.2 km²), of which 3,750 contain visible smoke plumes, accompanied by pixel-level masks.
@@
Reduced Inequalities (SDG-10): SustainFM includes an analysis of the sanitation index (unequal access to sanitary services), based on the DHS survey dataset from SustainBench [36]. The dataset utilizes a cluster-level toilet index to represent the sanitation index, ranked on a 1-5 scale, where 5 is the “highest quality”. The dataset comprises sanitation index for 89,271 clusters, computed from 2,143,329 households across 49 countries.
Sustainable Cities and Communities (SDG-11): SustainFM includes a building change detection task as a key proxy for urban dynamics by leveraging the LEVIRCD dataset [42]. The dataset captures significant urban development through Google Earth high-resolution imagery (0.5 m per pixel) across Texas, USA, from 2002 to 2018. It includes 637 image pairs (1024 × 1024 pixels), annotated with building additions and removals, encompassing diverse structures such as residential villas, high-rise apartments, garages, and warehouses.
@@
Responsible Consumption and Production (SDG-12) SustainFM addresses environmental degradation and conserving natural resources through monitoring changes caused by mining activities globally based on the MineNetCD dataset [43]. MineNetCD includes over 70,000 bitemporal image patches (256 × 256 pixels) from 100 global mining sites across six continents, collected via Google Earth Engine (GEE). Pixel-level change masks annotate mining expansion to support impact assessments and the pursuit of responsible production.
@@
Climate Action (SDG-13): The SustainFM benchmark includes a wildfire impact assessment task using the HLS Burn Scar Scenes dataset [44]. Based on the U.S. Monitoring Trends in Burn Severity (MTBS) database, this dataset combines burn scar shapefiles with Harmonized Landsat and Sentinel-2 (HLS) imagery acquired one to three months post-fire. It consists of 805 scenes, each measuring $512\times 512512 × 512$ pixels, with corresponding pixel-wise annotations of burn scars to support precise wildfire detection and post-event analysis.
Life below Water (SDG-14): SustainFM includes a task for detection of marine pollution—from oil slicks to plastic gyres—shedding light on eutrophication hotspots and floating debris through the Marine Debris and Oil Spills (MADOS) dataset [45]. This dataset comprises 174 Sentinel-2 scenes from 2015 to 2022, annotated with over 1.5 million pixels across diverse global waters and weather conditions. It captures pollution signatures at scale, spanning  47,000 m² of debris and  23 million m² of oil spills across 2803 patches ($240\times 240240 × 240$ px each, 2.4 km²).
@@
Life on Land (SDG-15): SustainFM includes a flood detection task to advance disaster response and land management built on the OMBRIA dataset [46]. This dataset integrates bitemporal, multimodal satellite data (Sentinel-1 SAR and Sentinel-2 multispectral imagery) paired with expert-labeled ground truth from ESA’s Copernicus Emergency Management Service. Covering 23 major flood events worldwide (2017–2021), OMBRIA comprises 3,376 annotated images ($256\times 256256 × 256$ px).
@@
Peace, Justice, and Strong Institutions (SDG-16): SustainFM contains a task to detect war-related damage using the GAZADeepDAV dataset [47], which features eight-band PlanetScope images at 3 m resolution. It comprises four pre- and four post-conflict scenes from Gaza in 2023, yielding 13,460 annotated patches ($256\times 256256 × 256$ px)—7,264 labeled as undamaged and 6,196 as damaged. GAZADeepDAV offers a unique foundation for developing EO tools that bolster transparency and resilience in conflict zones.
Evaluating Foundation Models on SustainFM Dataset
Preprocessing and Data Preparation
The SustainFM benchmark involves a variety of datasets with diverse spectral features, resolutions, and image sizes, which are designed for different AI applications, including segmentation, regression, and classification. This diversity brings difficulties to the evaluation of FMs on the SustainFM benchmark. Therefore, we preprocessed the dataset into a unified format to facilitate the data loading procedure to fine-tune and test the FMs. We resampled all the bands according to the highest resolution for the multispectral data, with different resolutions across the spectral bands. To ensure optimal performance and a convenient computation of the AI algorithms, we crop the satellite images of large sizes into small patches of around $256\times 256256 × 256$. To ensure data transparency and reproducibility, all preprocessed datasets have been uploaded to the Huggingface platform, which can be downloaded and reused in our implementation framework.
Foundation Models
Recent literature has witnessed rapid progress in developing FMs tailored to EO applications. We adopt the selection criteria proposed by PANGAEA [29] to ensure a rigorous and meaningful comparison, emphasizing reproducibility, methodological innovation, and scientific impact. Based on these criteria, we include 10 state-of-the-art FMs in our evaluation. Table 5 provides a comprehensive overview of the selected models, detailing their architectural scale, the datasets used for pretraining, and the associated EO sensors. These models represent diverse pretraining paradigms commonly employed in geospatial foundation modeling. Specifically, the selected FMs span three major pretraining strategies. Contrastive learning leverages auxiliary data in an unsupervised manner, aligning representations across different modalities to learn robust and transferable features. Masked image modeling employs a self-supervised approach, where portions of the input (i.e., patches of an image) are masked and the model is trained to reconstruct the missing content, thereby learning semantically meaningful representations. Supervised learning, by contrast, relies on labeled datasets to train the model directly for specific vision tasks, enabling the model to develop task-relevant perceptual capabilities. Below, we provide a detailed account of the ten FMs selected for evaluation, grouped according to their pretraining strategies: contrastive learning, masked image modeling, and supervised learning. These categories reflect the dominant paradigms currently shaping the development of general-purpose models in EO. An overview of the selected FMs is presented in Table 5, including their architectural backbones, pretraining datasets, learning strategies, model scales, and publication years.
  • Contrastive Learning: CROMA [18] aligns geographically and temporally co-registered SAR-optical image pairs using a contrastive objective, while also incorporating unimodal MIM losses to enhance modality-specific representations. RemoteCLIP [19] bridges the gap between vision and language by aligning satellite imagery with textual descriptions, resulting in visual representations enriched with high-level semantic concepts. SoftCon [20] introduces a multi-label supervision framework to learn cross-scene soft similarities, addressing challenges associated with rigid positive sample definitions and diverse semantic content in EO imagery.
  • Masked Image Modeling: DOFA [21] employs dynamic spectral adaptation to pretrain on multimodal EO data, explicitly incorporating physical characteristics of spectral bands based on wavelength information. GFM-Swin [22] applies cross-sensor reconstruction to learn generalizable representations across EO modalities. Prithvi [23] is the first FM pretrained on Harmonized Landsat and Sentinel-2 (HLS) data from NASA, integrating temporal dynamics during pretraining to capture seasonality and land cover evolution. SpectralGPT [10] adapts MIM for multispectral imagery, incorporating domain-specific mechanisms to accommodate spectral data’s unique structure and characteristics. ScaleMAE [24] addresses resolution heterogeneity in satellite data, learning robust representations from multimodal inputs at varying spatial scales. SSL4EO-S12 [25] serves as a foundational baseline FM, applying standard MIM techniques to Sentinel-1 and Sentinel-2 data in one of the earliest large-scale geospatial FM pretraining efforts.
  • Supervised Learning: SatlasNet [26] constructs a comprehensive multimodal dataset with diverse pretraining tasks, leveraging multi-task supervised learning to build an FM capable of capturing task-generalizable knowledge.
Table 5: Foundation models used in benchmarking, with architecture, pretraining data, strategy, size, and release year.
@@
Model Arch. Pretrained EO Data Learning Strategy Params (M) Year

### v2

@@ 2.1 No Poverty (SDG-1)
Therefore, SustainFM includes an asset wealth index regression task to assess localised poverty using the DHS data from SustainBench [43]. Wealth is estimated at the cluster level using a principal component analysis (PCA) of household assets, yielding a scalar index averaged across surveyed households in each cluster. The dataset comprises $2,079,036$ households across $86,936$ clusters in $48$ countries, based on DHS surveys conducted between 1996 and 2019.
2.2 Zero Hunger (SDG-2)
Ensuring food security and promoting sustainable agriculture is critical to combating hunger and malnutrition. Agricultural productivity is influenced by factors such as weather conditions, soil health, and water availability, all of which can be monitored using EO data. Satellite imagery enables large-scale assessment of cropland distribution, yield estimation, and the impact of natural disaster incidents [44, 45]. Within these perspectives, monitoring the fast and dynamic changes in cropland is a vital task to address food safety problems that are essential to zero hunger [46], which can provide timely cropland information to ensure cropland production and food security [47].
@@ 2.2 Zero Hunger (SDG-2)
SustainFM incorporates the Cropland Change Detection (CLCD) dataset to support food security and sustainable agricultural land management [48], as shown in Fig. 3. It comprises 600 cropland change samples derived from the Gaofen-2 satellite in Guangdong Province, China, in 2017 and 2019. Each group of samples consists of two images of $512\times 512$ pixels and a corresponding binary label indicating the cropland change. Each pair is annotated with binary labels indicating cropland conversion, including changes to buildings, roads, lakes, and bare soil, which are the key factors that threaten agricultural sustainability.
Figure 3: SDGs tasks and data samples in the SustainFM benchmark.
2.3 Good Health and Wellbeing (SDG-3)
Improving health outcomes and ensuring access to healthcare services are vital for enhancing global well-being. Health outcomes are shaped by a complex interplay of environmental and infrastructural factors, including air quality, climate, land use, and access to clean water and healthcare services. EO data offer spatially explicit indicators that can support public health interventions by identifying at-risk populations and monitoring environmental stressors associated with disease and mortality [49, 50, 51].
The SustainFM involves a task of children’s health regression based on the DHS survey dataset of SustainBench [43]. Specifically, the children’s health is indicated by the cluster-level child mortality rate that covers children who were age 5 or younger at the time of the DHS survey or who had died (age 5 or younger) no earlier than the year prior to the DHS survey. After identifying the qualifying children, the child mortality rate is further calculated as the number of deaths per $1,000$ children within each cluster. This dataset consists of $105,582$ cluster-level labels for child mortality rates computed from $1,936,904$ children under age 5.
2.4 Quality Education (SDG-4)
Access to quality education is essential for promoting equitable development and reducing social inequalities. Educational access is often influenced by geographical factors, including the proximity of schools, transportation networks, and infrastructure quality. Although EO data cannot directly measure educational outcomes, the triggers for quality education, such as school locations, population density, and education accessibility, are observable on satellite images [52, 53]. This information helps identify educational disparities, especially in rural and marginalized areas.
@@ 2.4 Quality Education (SDG-4)
Therefore, SustainFM includes an analysis of women’s educational attainment, based on the DHS survey dataset from SustainBench [43]. The educational attainment metric is derived by calculating the cluster-level mean of education years for women aged 15-49. To mitigate outliers, education years are capped at 18, a common threshold in surveys. Finally, the dataset includes $122,435$ cluster-level labels, computed from data on $3,013,286$ women across 56 countries.
2.5 Gender Equality (SDG-5)
Achieving gender equality requires eliminating discrimination and ensuring equal access to resources, services, and opportunities. EO-derived indicators such as vegetation cover, urbanization levels, and environmental quality can provide valuable perspectives into the underlying social, economic, and environmental conditions that influence women’s health [41]. By linking these spatial features to BMI outcomes, researchers and policymakers can identify regions where women are particularly vulnerable to malnutrition or obesity due to systemic inequality. This enables the development of geographically targeted, evidence-based interventions aimed at improving women’s health and well-being, thereby addressing one of the critical dimensions of gender inequality.
The SustainFM involves a task of women’s BMI index regression from satellite images based on the DHS dataset of SustainBench [43]. Predicting the BMI from the EO data has demonstrated effectiveness in several previous studies [54, 55], while the dataset from SustainBench solely focuses on the BMI indexes surveyed from women. In particular, this dataset contains $94,866$ cluster-level BMI labels computed from $1,781,403$ women of childbearing age (15-49), excluding pregnant women.
2.6 Clean Water (SDG-6)
Ensuring universal access to clean water and sanitation is vital for public health and sustainable development. Water scarcity, pollution, and uneven distribution of freshwater resources pose major challenges [56]. EO data plays a key role in monitoring surface water extent and quality [57, 58]. In the residential area, the black and odorous water (BOW) seriously affects the ecological balance of rivers and the health of people living nearby. Therefore, detecting the appearance of such water with EO data helps the authorities with water environment treatment that can improve the water quality in accurate, localized areas.
@@ 2.6 Clean Water (SDG-6)
SustainFM includes an urban black and odorous water (BOW) detection task using the RSBD dataset [59], which comprises 329 multispectral images from the Gaofen-2 satellite in Yantai, China, as shown in Fig. 3. The RSBD dataset is constructed in Yantai, a major port city in the Bohai Sea region in China, which is bordered by the Yellow Sea to the south and the Bohai Sea to the north. Since the northern coastal areas of China have experienced increasing water and energy shortages, as well as offshore environmental pollution, the BOW in the urban area has become a major obstacle to the development of Yantai City, which has limited the sustainable development of the regional economy and society. Moreover, pixel-level annotations of BOW enable the training of segmentation models to detect contaminated water, thereby supporting water quality management in urban settings.
2.7 Affordable and Clean Energy (SDG-7)
Reliable access to clean energy is critical for sustainable development and climate mitigation. With renewables surpassing fossil fuels in several countries (e.g., Germany, for instance, reached a 52% renewable share in 2024 [60]), tracking the deployment of clean energy infrastructure is increasingly important. EO data enables large-scale identification and monitoring of energy assets, such as solar farms, informing the global transition to low-carbon systems [61, 62, 63].
@@ 2.7 Affordable and Clean Energy (SDG-7)
SustainFM includes a solar farm identification task based on the GloSoFarID dataset [64], which provides $13,703$ multispectral Sentinel-2 image samples with pixel-level annotations. The dataset includes global solar farm samples from 2017 to 2018 and U.S. samples from 2022 to 2023, covering installations at a 10m spatial resolution (shown in Fig. 3). These EO-derived maps support the analysis of solar infrastructure expansion and guide decision-making in clean energy planning.
2.8 Decent Work and Economic Growth (SDG-8)
Promoting sustained economic growth and ensuring decent work opportunities are central to reducing poverty and inequalities. Economic development can be indicated by many factors such as urbanization, transportation networks, and infrastructural construction. In modern society, these indicators are significantly influenced by the electricity supply, which is fundamental to everyday life and industrial production [65]. However, some regions, especially in less developed countries, still lack an electricity supply, which not only significantly limits economic development but also affects the living quality of local residents [66, 67].
The SustainFM involves the task of detecting the settlements without electricity based on the detection of settlements without electricity (DSE) dataset [68]. This dataset was released as part of the 2021 data fusion contest organized by the IEEE Geoscience and Remote Sensing Society (GRSS). The dataset is derived from the Dezda and Salima sectors in Malawi, comprising 98 tiles of multimodal EO data collected from Sentinel-1, Sentinel-2, Landsat 8, and VIIRS nighttime satellites, as shown in Fig. 3. Each tile is of the size $800\times 800$ pixels, with each corresponding to a 64 $\mathrm{km}^{2}$ area. The EO data is then mapped into four classes, concerning the existence of human residents and the availability of the electricity supply.
2.9 Industry, Innovation, and Infrastructure (SDG-9)
Fostering industry, innovation, and infrastructure requires the development of sustainable, resilient systems that are responsive to environmental and societal challenges [69]. EO data can support these objectives by providing timely, spatially detailed insights into land use, urban growth, environmental conditions, and infrastructure health. Specifically, EO data enables smarter planning, risk assessment, and innovation in infrastructure design and industrial development, contributing to more sustainable and inclusive industrial growth. For example, detecting smoke plumes through satellite imagery can help monitor industrial emissions, offering valuable information for managing air quality and ensuring the continuity of critical infrastructure [70, 71, 72].
Therefore, the SustainFM dataset involves detecting the industrial smoke plumes from EO data based on the SmokePlume dataset [73]. The dataset acquires geographic locations of 624 sites from the European Pollutant Release and Transfer Register, which is a pollution reporting entity within the European Union. Subsequently, the dataset retrieved Sentinel-2 satellite imagery taken during 2019 for each site. Each raster image consists of all 12 spectral-band channels from the calibrated Level-2A reflectances and is cropped to a patch with a size of 120×120 pixels, corresponding to an area of $1.2\times 1.2\mathrm{km}^{2}$, as shown in Fig. 3. The dataset includes 21,350 images with 3,750 positive (a smoke plume is present) and 17,600 negative (no smoke plume is visible) samples. The dataset also generated masks indicating the pixels of smoke plumes for the positive images.
2.10 Reduced Inequalities (SDG-10)
Reducing inequality involves addressing income disparities, social exclusion, and unequal access to resources and services. Inequality often manifests spatially, with marginalized communities having poorer infrastructure and fewer services. Unequal access to sanitary services often leaves marginalized groups, such as low-income households, rural populations, and people living in informal settlements, without safe and dignified sanitation and threatens public health. EO provides an effective means to delineate these regions and analyze their socio-economic patterns, enabling the estimation of a continuous index that characterizes different degrees of deprivation [74].
@@ 2.10 Reduced Inequalities (SDG-10)
SustainFM includes an analysis of the sanitation index (unequal access to sanitary services), based on the DHS survey dataset from SustainBench [43]. The dataset utilizes a cluster-level toilet index to represent the sanitation index, ranked on a 151-5 scale, where $5$ is the “highest quality”. The dataset comprises a sanitation index for $89,271$ clusters, computed from $2,143,329$ households across 49 countries.
2.11 Sustainable Cities and Communities (SDG-11)
As urban populations surge, sustainable city development is crucial for managing infrastructure growth, mitigating environmental degradation, and ensuring equitable living conditions. Building construction serves as a key indicator of urban dynamics, reflecting shifts in population, economy, and land use. Monitoring these changes is essential for managing sustainable urban development. EO-based building change detection enables timely assessment of construction trends, informal settlements, and green space loss, informing urban planning, disaster response, and infrastructure equity [75, 76].
The SustainFM benchmark involves building change detection from remote sensing imagery based on the LEVIRCD dataset [77]. The LEVIRCD dataset focuses on significant land-use changes related to building, especially construction growth, spanning 5 to 14 years from 2002 to 2018 in the Texas region, U.S. It covers various types of buildings, such as villa residences, tall apartments, small garages and large warehouses. The dataset consists of $637$ very high-resolution (VHR, $0.5$ m/pixel) Google Earth (GE) image patch pairs, each with a size of $1024\times 1024$ pixels. Annotations of building additions and removals encompass a diverse range of structures, including residential villas, high-rise apartments, garages, and warehouses. This supports spatially explicit modeling of urban growth for sustainable city planning.
2.12 Responsible Consumption and Production (SDG-12)
Promoting sustainable consumption and production is critical for reducing environmental impacts and conserving resources. Unsustainable practices, such as overextraction, waste mismanagement, and pollution, contribute to environmental degradation, which the EO data can accurately capture. Specifically, land use and land cover changes can provide a timely footprint on human activities, enabling the tracking of their responsibility. Meanwhile, multitemporal EO data offers a powerful means to monitor human activities, allowing for the dynamic tracking of land use changes linked to resource exploitation [78, 79, 80].
@@ 2.12 Responsible Consumption and Production (SDG-12)
SustainFM addresses environmental degradation and conserves natural resources by monitoring changes caused by mining activities globally, based on the MineNetCD dataset [81]. As shown in Fig. 3, MineNetCD includes over $70,000$ bitemporal image patches ($256\times 256$ pixels) from 100 global mining sites across six continents, collected via Google Earth Engine (GEE). Pixel-level change masks facilitate the assessment of mining expansion, supporting impact assessments and the pursuit of responsible production.
2.13 Climate Action (SDG-13)
Addressing climate change requires reducing greenhouse gas emissions, mitigating environmental degradation, and enhancing resilience to extreme weather events. EO data plays a crucial role in tracking climate variables, including land surface temperature, deforestation, and sea-level rise, as well as monitoring hazards such as floods, droughts, and wildfires [82]. In particular, mapping wildfires from satellite imagery provides spatial insight into fire frequency, intensity, and burn extent, informing carbon emission estimates and ecosystem recovery efforts [83, 84]. These data-driven insights support disaster preparedness and enable adaptive responses to climate-related risks.
@@ 2.13 Climate Action (SDG-13)
The SustainFM benchmark includes a wildfire impact assessment task using the HLS Burn Scar Scenes dataset [85], as shown in Fig. 3. Based on the U.S. Monitoring Trends in Burn Severity (MTBS) database, this dataset combines burn scar shapefiles with Harmonized Landsat and Sentinel-2 (HLS) imagery acquired one to three months post-fire. It is created based on the Monitoring Trends in Burn Severity (MTBS) historical fire database, constructed in the U.S. region, which provides the location of wildfire emergence and the shape file of burn scars. It consists of 805 scenes, each measuring $512\times 512$ pixels, with corresponding pixel-wise annotations of burn scars to support precise wildfire detection and post-event analysis.
2.14 Life below Water (SDG-14)
Safeguarding marine ecosystems requires urgent action to mitigate pollution and reduce anthropogenic stressors. EO provides a critical means of monitoring ocean health, enabling systematic assessment of sea surface temperature, chlorophyll concentration, turbidity, and harmful algal blooms. Moreover, EO can identify marine pollution ranging from oil slicks to plastic gyres, revealing eutrophication hotspots and floating debris to support early warning systems and targeted interventions [86, 87].
The SustainFM benchmark involves a task of detecting the marine debris and oil spill in the ocean based on the MADOS dataset [88], as shown in Fig. 3. It is composed of multispectral Sentinel-2 data, consisting of 174 scenes captured between 2015 and 2022, with approximately 1.5M annotated pixels, which are globally distributed and collected under various weather conditions. The dataset focuses on a wide range of classes of marine objects related to the pollutants in the form of marine debris and oil spills, which cover $4696$ and $234,568$ pixels. The dataset contains 2803 annotated image patches with a size of $240\times 240$, each covering an area of $2.4\times 2.4\mathrm{km}^{2}$.
2.15 Life on Land (SDG-15)
Protecting terrestrial ecosystems hinges on addressing threats from both natural and human activities. EO data is indispensable for tracking landscape dynamics, from shifts in vegetation cover and soil moisture to the aftermath of natural disasters [89]. Flood mapping via EO, for instance, enables rapid assessment of inundation impacts on forests, croplands, and habitats, which are vital for risk mitigation, restoration, and land stewardship [90, 91, 92].
@@ 2.15 Life on Land (SDG-15)
SustainFM includes a flood detection task to advance disaster response and land management, which is built on the OMBRIA dataset [93], as shown in Fig. 3. This dataset integrates bitemporal, multimodal satellite data derived from Sentinel-1 SAR and Sentinel-2 multispectral imagery—paired with expert-labeled ground truth from ESA’s Copernicus Emergency Management Service. Covering 23 major flood events worldwide (2017–2021), OMBRIA comprises 3,376 annotated images ($256\times 256$ px).
2.16 Peace, Justice, and Strong Institutions (SDG-16)
Promoting peace, justice, and strong institutions requires protecting civilians, upholding human rights, and ensuring accountability during and after conflict. EO data contributes to these goals by enabling the classification and monitoring of war-induced damage to infrastructure, settlements, and cultural heritage sites [94]. High-resolution satellite imagery can detect changes such as building destruction, displaced populations, and disrupted land use, providing objective evidence in near real-time [95]. This information supports humanitarian response, aids in post-conflict reconstruction, and can serve as documentation for legal and human rights investigations.
@@ 2.16 Peace, Justice, and Strong Institutions (SDG-16)
SustainFM contains a task to detect war-related damage using the GAZADeepDAV dataset [96], which features eight-band PlanetScope images at 33m resolution (shown in Fig. 3. It comprises four pre- and four post-conflict scenes from Gaza in 2023, yielding $13,460$ annotated patches ($256\times 256$ px), from which $7,264$ were labeled as undamaged and $6,196$ as damaged. GAZADeepDAV provides a distinctive foundation for developing EO tools that enhance transparency and resilience in conflict zones.
Table 2: Overview of the foundation models for benchmarking.
@@
Model Architecture Pretrained EO Data Learning Strategies Parameters (M) Year